h 5. In Homework 8 we looked at the finite difference approximation of f(x). f(x + 1) = f(x) + f(z)h + }} (-x)+2 +0(h2). R
Answer #1

His Solution f(x) = Cos (710) x=13 h=0.5 =) h²=0.25 h3 = 0.125 f(x) = -K sin (AX) f(x) = -72 Cos (TX) f(x+h) = f(x) + f(x+